# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v2

@@ 4 Fixing the identifiability problem: the Hypergeometric correlation model
In this section we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints. We then address the identifiability issue discussed in the previous section by introducing a parsimonious subclass selected according to an explicit optimality criterion.
@@
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l.$ (11)
@@
The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
@@
Let $d$ be fixed, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$, and $l\geq 0$. Then:
@@
    and only if
    $\mu\geq\delta-l+\frac{1}{2}.$
For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered as a special case. Specifically, since $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ if and only if $\kappa\geq\frac{1}{2}-\frac{d}{2}$, which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, or for $d=1$ and $\kappa\geq 0$, Regime 1 of Theorem 2 gives the necessary and sufficient condition
$\mu\geq\frac{d+1}{2}+\kappa,$
For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq 0.5(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
As discussed in Section 3, the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under the interchange $\beta\leftrightarrow\gamma$, corresponding to $l\mapsto-l$ in the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. The admissible region is therefore symmetric with respect to $l=0$, and restricting attention to $l\geq 0$ entails no loss of generality.
Translating Theorem 2 back to the original parameterization $\mathcal{GH}_{\delta,\beta,\gamma,a}$, the following equivalent conditions hold for $\delta>d/2$ and $a>0$:
@@
    and only if
    $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
In particular, in Regime 1 ($0\leq\gamma-\beta\leq\delta-\frac{1}{2}$), condition (c) of Section 3 implies condition (b), so that (a) and (c) alone characterize validity. Hence, within this regime, the sufficient conditions of Emery and Alegría (2022) reduce to (a) and (c) and are therefore also necessary. In Regime 2 ($\gamma-\beta>\delta-\frac{1}{2}$), whether the gap between sufficiency and necessity can be closed remains an open question.
We now turn to the structural identifiability issue for the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model. To resolve it, the parameter $l$ must either be fixed (as in the $\cal GW$} model) or linked to the remaining parameters and/or to the spatial dimension. This naturally raises the question: which choice of $l\geq 0$ is optimal ?
We address this question by introducing an explicit optimality criterion that selects a single parsimonious representative once smoothness and compact support are fixed. We adopt the criterion of maximum integral range attainable by a compactly supported correlation with given smoothness and support.
@@
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
This quantity, originally introduced by Lantuéjoul (1991), provides a global and threshold-free measure of dependence (Yaglom, 1987; Matheron, 1989; Hristopulos and Zukovic, 2011). If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$, so that the integral range coincides with the spectral density at zero frequency. When $d=1$, AA reduces to the area under the correlation function.
For compactly supported correlations, maximizing $A$ therefore selects, among all admissible models with fixed smoothness and support, the one with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints. While alternative optimality principles could be considered, they would typically depend on the sampling design or on a specific loss function. In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification within the family.
@@
$A_{\nu,\alpha,d}=\frac{\alpha^{d}\pi^{d/2}\Gamma\!\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (13)
@@
The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by:
$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+\frac{1}{2})\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (14)
Note that, as expected, the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation. In addition, using asymptotic properties of the Gamma function, it can be shown that $A_{\kappa,\mu,l,a,d}\sim\mu^{-d}$ as $\mu\to\infty$ and, as a consequence,
@@
This implies that, as the shape parameter $\mu$ increases, the integral range approaches zero, regardless of the compact support and smoothness parameters. The following theorem establishes that, for a given $l\geq 0$, the maximum integral range is obtained by fixing $\mu$ on the lower bound of its parametric space.
@@
Let $\kappa>0$, $a>0$, and $d\in\mathbb{N}$ be fixed. For each $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ with respect to $\mu$ (over its admissible parameter space) is attained at $\mu^{\star}(l)$, where:
@@
    if $0\leq l\leq\dfrac{d}{2}+\kappa$, then
    $\mu^{\star}(l)=\mu_{1}(l):=\delta-l+\frac{1}{2};$
@@
    if $l>\dfrac{d}{2}+\kappa$, then
    $\mu^{\star}(l)=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l.$
For instance, let $\kappa\geq 0$. If $l=\frac{1}{2}$, then from (9) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of $\cal GW$} model in this case.
As an example, Figure 1 (left part) depicts the ${\cal GW}_{0,\mu,1}$ model and the right part depicts the ${\cal GW}_{1,\mu,1}$ model for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$. In these examples $d=1$ and it is apparent that the model with maximum area corresponds to the case $\mu=\mu_{1}(\frac{1}{2})$ (red line) in both examples.
The next step is to find $l\geq 0$ that maximizes $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$.
@@
Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d}{2}+\kappa$.
@@
    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,\delta]$, hence
    $\max_{0\leq l\leq\delta}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,\mu_{1}(\delta),\delta,a,d}=A_{\kappa,1,\delta,a,d}.$
@@
    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(\delta,\infty)$ and
    $\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,\delta,a,d}.$
In particular, the global maximum over all admissible $(\mu,l)$ is attained at $(\mu,l)=(1,\delta)$ and is equal to $A_{\kappa,1,\delta,a,d}$.
@@
Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$. Right part: the same Figure but with $\kappa=1$. In both cases $d=1$.
As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines, respectively). In particular the left part depicts the case $\kappa=0$ and the right part depicts the case $\kappa=1$ ($d=2$ in both cases). It is apparent that the case $l=\frac{d}{2}+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the $\cal GW$} models (red line).
@@
Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines respectively). Right part: the same Figure but with $\kappa=1$. The case $l=\frac{1}{2}$ (red line) corresponds to the $\mathcal{GW}_{\kappa,\mu,1}$ model. The blue line corresponds to the correlation model that maximizes the integral range. In both cases $d=2$.
Summarizing our results, the model that can potentially achieve the maximum integral range is $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ and, in particular, the maximum is achieved when $\mu=\mu_{1}(d/2+\kappa)=1$. Taking into account (15), this implies that
$0\leq A_{\kappa,\mu,l,a,d}\leq a^{d}\left[\frac{\pi^{d/2}\,\Gamma\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\left(1+\kappa\right)\,\Gamma\left(1+\frac{d}{2}+2\kappa\right)}{2^{d}\,\Gamma\left(\kappa+\frac{1}{2}\right)\,\Gamma\left(1+\frac{d}{2}+\kappa\right)\,\Gamma\left(1+d+2\kappa\right)}\right].$
Accounting for (11), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron (1965, I.4.18).
Fixing $l=\frac{d}{2}+\kappa$, we define the hypergeometric correlation by
@@
Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized. Indeed, by setting $l=\frac{d}{2}+\kappa$ in Theorem 2, we obtain the condition
@@
The next theorem shows that the subclass ${\cal H}_{\kappa,1,a,d}$, obtained by maximizing the integral range, retains substantial modeling flexibility. Through beta-type scale mixtures, it can generate any member of ${\cal H}_{\kappa,\mu,a,d}$ and a broad subclass of the Gauss–hypergeometric family $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. Thus, the maximal integral range criterion does not restrict modeling capacity, but instead selects a parsimonious representative that remains rich within the admissible class.
Theorem 6.
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$, and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ by Theorem 2. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of hypergeometric correlations $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
In particular, when $l=\frac{d}{2}+\kappa$, one obtains
Clearly, the ${\cal H}_{\kappa,\mu,a,d}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$ in the same way as the ${\cal GW}_{\kappa,\mu,\beta}$ model. Figure 3 depicts two realizations of zero-mean and unit-variance Gaussian RFs with correlations ${\cal H}_{\kappa,4,0.2,2}$ for $\kappa=0,1,2$. The realizations were constructed via the Cholesky decomposition of the covariance matrix.

### v3

@@ 4 Fixing the identifiability problem: the Hypergeometric correlation model
In this section, we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints. We then address the non-uniqueness discussed in the previous section by imposing an appropriate restriction on the parameter space, which restores a unique representation of the correlation function. Finally, we introduce a parsimonious compactly supported subclass selected through an explicit optimality criterion.
@@
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l,\qquad l\geq 0.$ (11)
@@
The restriction $l\geq 0$ in (11) is equivalent to imposing the ordering constraint $\gamma\geq\beta$ on the original $\mathcal{GH}_{\delta,\beta,\gamma,a}$ family. This removes the intrinsic label-switching ambiguity induced by the symmetry of $_{2}F_{1}$ in its upper parameters, namely $\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a}$, and thus yields a unique representation up to this exchange.
The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for the reparameterized model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
@@
Let $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$. Then:
@@
    and only if $\mu\geq\delta-l+\frac{1}{2}.$
Remark 1.
Tracing back the conditions of Theorem 2 to the original parametrization $(\delta,\beta,\gamma,a)$, and assuming $\delta>d/2$ and $\gamma\geq\beta$ we obtain:
@@
    and only if $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
In particular, on the broad regime $\gamma-\beta\leq\delta-\tfrac{1}{2}$ the sufficient conditions (a)–(c) in Section 3 reduce to the single inequality $\beta+\gamma\geq 3\delta+\tfrac{1}{2}$ and become sharp, since this inequality is also necessary. This sharpens the earlier validity region reported in Emery and Alegría (2022). Outside this regime, the sufficient bound above is explicit while a residual gap with the necessary condition remains.
For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered. Indeed, if $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ (which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, and for $d=1$ with $\kappa\geq 0$), then Regime 1 of Theorem 2 yields $\mu\geq\delta=\frac{d+1}{2}+\kappa$. For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq\frac{1}{2}(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
Theorem 2 characterizes validity for each admissible value of $l$. However, even after restricting to $l\geq 0$ the reparameterized model still involves two shape parameters, $\mu$ and $l$. In the spirit of the $\cal GW$} model, we therefore seek a parsimonious specification by treating $l$ as fixed (or by linking it to the remaining parameters and/or to the spatial dimension), so that inference targets a single interpretable shape parameter $\mu$. This naturally raises the question: which choice of $l\geq 0$ is optimal? We address this question by introducing an explicit optimality criterion that selects a parsimonious subclass once smoothness and compact support are fixed. Specifically, we choose $l$ so that the resulting $\mathcal{GH}$ subclass can attain the largest isotropic integral range.
@@
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
This quantity, originally introduced by Lantuéjoul (1991), provides a global and threshold-free measure of dependence (Yaglom, 1987; Matheron, 1989; Hristopulos and Zukovic, 2011).
If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then
$A=(2\pi)^{d}\,\widehat{\phi}(0),$
so the integral range is proportional to the spectral density at zero frequency.
For compactly supported correlations, maximizing $A$ therefore targets models with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints. While alternative optimality principles could be considered, they would typically depend on the sampling design or on a specific loss function. In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification (through $l$) within the family.
@@
$A_{\nu,a,d}=2^{d}\pi^{d/2}a^{d}\,\frac{\Gamma\!\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (13)
@@
The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by
$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\,\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{\Gamma(\kappa+\frac{1}{2})\,\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\,\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (14)
Note that the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation. In addition, using asymptotic properties of the Gamma function, it can be shown that $A_{\kappa,\mu,l,a,d}\sim\mu^{-d}$ as $\mu\to\infty$ and, as a consequence,
@@
This implies that, as the shape parameter $\mu$ increases, the integral range approaches zero, regardless of the compact support and smoothness parameters.
The following theorem establishes that, for a given $l\geq 0$, the integral range is maximized by taking $\mu$ at the lower bound of its admissible domain. In what follows, we call admissible the parameter region ensured by Theorem 2. In Regime 2 of Theorem 2, validity is guaranteed only through the explicit sufficient bound. Throughout Sections 45, whenever we optimize over $(\mu,l)$ we therefore interpret the admissible region as the parameter set ensured by the sufficient conditions in Theorem 2.
@@
Let $\kappa>-\frac{1}{2}$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d+1}{2}+\kappa$. For each $l\geq 0$, the map $\mu\mapsto A_{\kappa,\mu,l,a,d}$ is non-increasing on its admissible domain. Hence its maximum is attained at the smallest admissible value of $\mu$. In particular,
@@
    if $0\leq l\leq d/2+\kappa$, then the maximum is attained at μ^(l)=μ_1(l):=δ-l+12;
@@
    if $l>d/2+\kappa$, then the maximum is attained at μ^(l)=μ_2(l):=2κ+l^2+d+1-l.
For instance, if $l=\frac{1}{2}$ and $d\geq 2$ (or $d=1$ and $\kappa\geq 0$), then by (9)
$\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a},$
and the maximum integral range (over the admissible $\mu$) is attained at $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which coincides with the lower bound of the $\cal GW$} model in this case.
As an example, consider the case $d=1$ and $\kappa=0,1$. Figure 1 (left) depicts ${\cal GW}_{0,\mu,1}$ and Figure 1 (right) depicts ${\cal GW}_{1,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$. It is clear that, in both panels, the correlation with the largest area under the curve is obtained for $\mu=\mu_{1}\!\left(\tfrac{1}{2}\right)$ (red line).
The next step is to determine which $l\geq 0$ yields the largest attainable integral range, i.e., which $l$ maximizes $A_{\kappa,\mu^{\star}(l),l,a,d}$.
@@
Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed.
@@
    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,d/2+\kappa]$, hence
    $\max_{0\leq l\leq d/2+\kappa}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,1,d/2+\kappa,a,d}.$
@@
    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(d/2+\kappa,\infty)$ and
    $\lim_{l\downarrow d/2+\kappa}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,d/2+\kappa,a,d}.$
In particular, the largest attainable integral range over all admissible $(\mu,l)$ is achieved at $(\mu,l)=(1,d/2+\kappa)$ and equals $A_{\kappa,1,d/2+\kappa,a,d}$.
@@
Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$. Right part: the same Figure but with $\kappa=1$. In both cases $d=1$.
As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},d/2+\kappa$ (black, red and blue lines, respectively). In particular the left part depicts the case $\kappa=0$ and the right part depicts the case $\kappa=1$ ($d=2$ in both cases). It is apparent that the case $l=d/2+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the $\cal GW$} models (red line).
@@
Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},d/2+\kappa$ (black, red and blue lines respectively). Right part: the same Figure but with $\kappa=1$. The case $l=\frac{1}{2}$ (red line) corresponds to the $\mathcal{GW}_{\kappa,\mu,1}$ model. The blue line corresponds to the correlation model that maximizes the integral range. In both cases $d=2$.
Summarizing, Theorem 5 shows that the largest attainable integral range over all admissible $(\mu,l)$ is achieved at $(\mu,l)=(1,d/2+\kappa).$ In particular, fixing $l=d/2+\kappa$ identifies the unique $\mathcal{GH}$ specification that can attain this maximum (with the maximum attained at $\mu=1$).
Accounting for (11), the extremal member attaining the maximal integral range (i.e., $l=d/2+\kappa$ and $\mu=1$) is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$. This coincides with the Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular/tent for $d=1$, circular for $d=2$, spherical for $d=3$, pentaspherical for $d=5$, see Section 5), and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained via a (fractional) montée (Matheron, 1965, I.4.18).
Motivated by this, we fix $l=d/2+\kappa$ and define the hypergeometric correlation by
@@
Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized. Indeed, by setting $l=d/2+\kappa$ in Theorem 2, we obtain
@@
The next theorem shows that the extremal subclass ${\cal H}_{\kappa,1,a,d}$, which attains the maximal integral range, retains substantial modeling flexibility. Through beta-type scale mixtures, it can generate any member of ${\cal H}_{\kappa,\mu,a,d}$ and a broad subclass of the Gauss–hypergeometric family $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. Thus, the maximal integral range criterion does not restrict modeling capacity, but instead selects a parsimonious representative that remains rich within the admissible class.
Theorem 6.
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$ and $l\geq 0$. Assume that the parameters $(\mu,l)$ belong to the admissible region of Theorem 2, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$. In addition, assume $\mu+2l>1+d+2\kappa$. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of $\mathcal{H}_{\kappa,1,b,d}$ with $b=a\sqrt{uv}$:
In particular, when $l=d/2+\kappa$, one obtains
[…]